A source vibration creates waves through a medium with a wavelength of 43.0 m.
If the speed of the wave is 520 km/s, then:
a) what is the source frequency?
b) how much time would it take this wave to travel 800 m?
1
Expert's answer
2020-06-16T09:29:42-0400

a) A wavelength "\lambda" and a speed of traveling wave "v" are connected with the frequency "f" in the following manner:


"f = \dfrac{v}{\lambda} = \dfrac{43cm}{520km/s}= \dfrac{0.43m}{520\cdot 10^3m/s} \approx 0.83\mu Hz"

b) As far as the speed of the wave is given, the time to cover the distance 800m can be found as following:


"t = \dfrac{800m}{v} = \dfrac{0.8km}{520km/s} \approx 1.5ms"

Answer. a) f = 0.83 uHz, b) t = 1.5 ms.